OpenStudy (dtan5457):

Still not getting these rational expressions 100%...

OpenStudy (dtan5457):

\[(3y^-2+y^3)^-1\]

OpenStudy (dtan5457):

those are all powers

TheSmartOne (thesmartone):

\[(3y^{-2}+y^3)^{-1}\]

TheSmartOne (thesmartone):

Like that?

OpenStudy (dtan5457):

Yes. Idk how to make negative exponents..

TheSmartOne (thesmartone):

After ^ this you put { and } with the negative inside too

TheSmartOne (thesmartone):

\(\Huge x^{-1}=\frac{1}{x}\)

OpenStudy (dtan5457):

yes, i put it all under the denominator

TheSmartOne (thesmartone):

Just use that and put the rational expression at the denominator and its solved...

OpenStudy (dtan5457):

for some reason, the numerator has a y^2 has the answer..

OpenStudy (dtan5457):

where does that come from?

TheSmartOne (thesmartone):

must be a typo. :/

OpenStudy (dtan5457):

wow are you serious? stupid barrons book that would be the 2nd typo in the answer and it's only page 42..

TheSmartOne (thesmartone):



OpenStudy (dtan5457):

non-like bases, cant add them , forgot

TheSmartOne (thesmartone):

Hmm. I din't think so. @Nnesha What do you think?

OpenStudy (dtan5457):

the apparent answer is \[\frac{ y^2 }{ 3+y^5 }\]

OpenStudy (dtan5457):

that's the first answer i got, was sure of it, only to see it's wrong..

OpenStudy (mathmath333):

\(\Huge \checkmark\)

OpenStudy (mathmath333):

\[\frac{ y^2 }{ 3+y^5 }\]